# Who Gets Cited Most? Benchmarking Long-Context Reasoning on Scientific Articles

### v2

@@ Abstract
We introduce SciTrek, a novel question-answering benchmark designed to evaluate long-context reasoning capabilities of large language models (LLMs) using scientific articles. Current long-context benchmarks often focus on simple information retrieval tasks, or employ artificial contexts. SciTrek addresses these limitations by creating benchmark questions that require information aggregation and synthesis across multiple full-text scientific articles. The questions and their ground-truth answers are automatically generated by formulating them as SQL queries over a database constructed from article metadata (i.e., titles, authors, and references). These SQL queries provide explicit, verifiable reasoning processes that enable fine-grained error analysis on model answers, and the data construction scales to contexts of up to 1M tokens with minimal supervision. Experiments on open-weight and proprietary LLMs show that SciTrek poses significant challenges as the context length increases, with supervised fine-tuning and reinforcement learning offering only limited gains. Our analysis reveals systematic shortcomings of frontier LLMs’ ability to effectively perform numerical operations and accurately locate information in long contexts. [^1] Our dataset and code are available from https://github.com/oaimli/SciTrek
@@ 1 Introduction
Large language models (LLMs) show promise in accelerating scientific progress by assisting researchers in reviewing and synthesizing the growing body of literature (Pearson, 2024; Agarwal et al., 2025) and generating novel research ideas (Si et al., 2025; Baek et al., 2025). So much so that industry products such as Deep Research [^2] https://openai.com/index/introducing-deep-research/, Elicit [^3] https://elicit.com/, and Scite [^4] https://scite.ai/ have emerged as dedicated tools to aid users with complex, multi-step research tasks. Scientific workflows often require processing large inputs: full scientific articles, collections of references, or structured datasets. A model capable of processing 100K+ tokens could simultaneously analyze multiple articles, track long chains of reasoning, or connect experimental results to prior work.
While long-context language models (LCLMs) are highly relevant to scientific tasks, existing benchmarks for evaluating their capabilities predominantly focus on non-scientific texts (Kuratov et al., 2024; Maekawa et al., 2025; Bai et al., 2025; Yen et al., 2025). They mostly address simple information retrieval tasks (e.g., Needle-In-A-Haystack; Kamradt 2023; Hsieh et al. 2024) rather than information integration, i.e., reasoning across multiple documents to synthesize information and answer complex questions. Some question-answering benchmarks (Asai et al., 2024; An et al., 2024; Bai et al., 2025) require information integration but do not provide a typology of the reasoning steps involved, making it difficult to determine which capabilities are being assessed and why a model fails.
Another issue relates to scalability and extensibility. When substantial human effort is required to generate benchmark questions and their answers (e.g., Asai et al. 2024), scaling to longer contexts or different types of questions becomes prohibitive. To better emulate scientific research practice, it is also important to evaluate a range of skills in natural contexts, such as a model’s ability to aggregate or filter information, and combinations thereof (Wang et al., 2024c). In contrast, a few benchmarks rely on artificial contexts, created by adding noise  (Kuratov et al., 2024) to short contexts (e.g., HotpotQA; Yang et al. 2018) or by synthesizing the input with LLMs (Maekawa et al., 2025).
In this paper, we create a question-answering benchmark based on scientific articles that aims to alleviate these issues. We assume a scenario where a researcher, possibly during the literature review process, seeks answers to specific questions when the related scientific articles are provided. As an example, consider the first question in Table 1: “What is the highest number of authors that any single article has?”. Although seemingly simple, it requires aggregation of information across multiple articles. To answer correctly, a hypothetical model would need to count the authors of each article in the collection and select the one with the largest number. We generate questions requiring explicit reasoning by focusing on basic numerical operations such as counting, aggregating, and sorting, applied to core elements of scientific articles: titles, authors, and references. Since these operations correspond to standard database functions, we can automatically generate arbitrarily complex SQL queries over any collection of scientific articles and obtain ground-truth answers without manual annotation by constructing database tables for titles, authors, and references. We exploit recent advances in LLM capabilities for SQL understanding and generation (Yu et al., 2018; Li et al., 2023; Hui et al., 2024) to automatically convert SQL queries into natural-language questions.
Our benchmark, SciTrek, contains a test set of 2,121 question-answer pairs on scientific articles and a training set of 19,543 question-answer pairs, with contexts of varying lengths (i.e., 64K, 128K, 512K, and 1M). [^5] In the paper, K refers to 1,024 and M to 1,024K. Table 1 shows examples of questions, their corresponding SQL queries, and answers. While these questions are relatively superficial and do not engage deeply with the domain-specific content, they require models to retrieve and synthesize facts that are sparsely distributed across article collections and span long contexts. If models struggle with simple comparison, filtering, sorting, and aggregation operations, they might be impossible to succeed on more complex analytical tasks (Wolfson et al., 2020, 2025). We use SciTrek to evaluate the reasoning capabilities of LCLMs with context windows exceeding 128K tokens. Experimental results reveal that SciTrek poses significant challenges for these LCLMs. In general, models achieve poor performance, even when enhanced with supervised fine-tuning (SFT) and reinforcement learning (RL). This paper makes the following contributions:
  • We propose a methodology for constructing a long-context question-answering benchmark over natural texts, where explicit reasoning processes are approximated via SQL operations involving information aggregation. The methodology requires minimal human intervention and generalizes to any domain with entities that can be structured as database tables.
  • Extensive experiments reveal that frontier open-weight and proprietary LLMs struggle with this task, especially as input length increases. While SFT and RL improve over baseline models, performance remains limited. All models perform poorly on questions involving references and citation relationships among articles.
  • By design, SciTrek enables fine-grained analysis of model behavior, offering insight into where and why models fail. Evaluation of open-weight models reveals systematic shortcomings in counting bibliographic elements and performing basic numeric operations, with error rates increasing as question length grows; models also frequently misinterpret compound conditions and struggle with logical constructs involving negation (see last question in Table 1).
@@
Dataset Natural Context #SciQs Reasoning Scalable Len     Example Question
@@
OpenScholar 208 3K Compile a list of reviews […], and identify the most promising […]
LongBench v2 50 128K How long have I been living in my current apartment in Shinjuku?
@@
Loong 53 250K Which company has the highest non-current assets?
SciTrek 1,716 1M How many articles are cited by other articles but do not cite any other articles?
Table 2: Representative benchmarks for evaluating LCLMs. Natural Context: does the benchmark have natural input contexts; #SciQs: the number of unique benchmark questions or instructions with ground-truth answers on scientific articles (—: benchmarks that are not using any scientific articles); Reasoning: does the benchmark provide detailed reasoning skills required to answer each question; Scalable: can the benchmark be extended with minimal effort (e.g., to longer contexts or larger datasets); Len: the maximum input length that the model supports in terms of tokens; : length based on texts retrieved rather than full input (which is not available).
@@ 2 Related Work
Long-context Language Modeling
A growing body of work explores the challenges LLMs encounter when processing inputs that exceed their context lengths. Rotary Position Embeddings (RoPE; Su et al. 2024) have emerged as a widely used approach that replaces absolute positional embeddings with rotational transformations of token embeddings, improving generalization to unseen sequence lengths. Building on RoPE, several extensions introduce rescaling to further enhance performance on substantially longer inputs without the need for retraining (Peng et al., 2024).
Other architectural advances focus on reducing the quadratic cost of the Transformer attention mechanism (Sun et al., 2025). Sparse attention methods, such as LongFormer (Beltagy et al., 2020), compute only selected portions of the full attention matrix, thereby reducing computational cost and accelerating both prefilling and inference (Jiang et al., 2024; Fu et al., 2024). Many contemporary models interleave sparse and full attention layers to balance efficiency and performance (Dubey et al., 2024; Llama Team, 2025; Yang et al., 2025; Kamath et al., 2025). Another popular strategy for reducing computation in long-context settings is to limit the number of active model parameters. Group Query Attention (GQA; Ainslie et al. 2023) enables multiple attention heads to share key–value projections, without significant performance loss. The open-weight LLMs used in our experiments combine these techniques, complemented by pretraining on long-context tasks (Dubey et al., 2024; Kamath et al., 2025; Yang et al., 2025; Llama Team, 2025).
Long-context Datasets
@@
The growing interest in LCLMs has driven the development of numerous benchmarks designed to evaluate their performance. Initial benchmarks primarily assessed whether models can retrieve information from their context, typically through Needle-In-A-Haystack (NIAH) tasks, such as NeedleBench (Li et al., 2024). Building on these approaches, subsequent benchmarks either added distractor information to short-context tasks or scattered relevant information across synthetic contexts, such as Ada-LEval (Wang et al., 2024a), BABILong (Kuratov et al., 2024), HELMET (Yen et al., 2025), LIFBENCH (Wu et al., 2025b), LongBench (Bai et al., 2024), and RULER (Hsieh et al., 2024).
In addition to retrieval, some benchmarks emphasize complex reasoning over long contexts. For example, OpenScholar (Asai et al., 2024) and LongBench v2 (Bai et al., 2025) contain questions that require extensive use of multiple documents provided as context, while DocFinQA (Reddy et al., 2024) and MedOdyssey (Fan et al., 2025) focus on financial and medical reasoning, respectively. Although not explicitly technical, LongMemEval (Wu et al., 2025a) and L-Eval (An et al., 2024) also require reasoning over natural, long-context information. However, evaluating model answers against these benchmarks remains challenging, as the reasoning steps required to arrive at correct responses are often opaque, making it difficult to assess failure modes. In contrast, HoloBench (Maekawa et al., 2025) and MathHay (Wang et al., 2024b) better delineate between the types of reasoning required for each question, but rely on unnatural contexts. CURIE (Cui et al., 2025), OpenScholar, and LongBench v2 evaluate reasoning over scientific articles, but their reliance on expert-curated questions and answers poses challenges for scalability to longer contexts or larger datasets.
In designing SciTrek, we aim to retain many of the useful qualities of previous benchmarks (e.g., structured reasoning, scalability) while still proposing a fairly natural and easy task for humans to perform. We compare SciTrek with existing benchmarks in Table 2. Unlike Loong (Wang et al., 2024c), which tests broad long-context comprehension, SciTrek uniquely targets scientific synthesis and structured reasoning at extreme context lengths. Although we focus on scientific articles, our methodology could extend to other domains with explicit entities and relations. Thanks to the SQL backbone, we are able to construct questions testing various model skills (e.g.,  aggregation, filtering), and how these manifest themselves through explicit reasoning processes.
@@ 3 The SciTrek Benchmark
Figure 1: Overview of SciTrek construction process: we gather article collections of varying scales; we then obtain SQL queries and their answers based on databases (which store collection-specific metadata); and finally we convert SQL queries to natural language questions. The final dataset consists of input contexts (full-text scientific articles), natural language questions, and their answers, all highlighted in blue.
Figure 1 illustrates how SciTrek was curated: we first gather scientific article collections as contexts corresponding to different lengths (i.e., 64K, 128K, 512K, and 1M); we then create databases representing the article metadata and SQL queries with answers from database execution; finally, we convert the SQL queries into natural language questions. We explain these stages in detail below.
@@ 3.1 Gathering Scientific Article Collections
SciTrek is constructed using scientific articles from Semantic Scholar. [^6] https://www.semanticscholar.org/ To cover diverse topics of articles, we obtain an initial set of seed articles from eight subjects: Computer Science (CS), Economics, Electronic Engineering (EE), Math, Physics, Biology, Finance, and Statistics. [^7] Following the typology of subjects from https://arxiv.org/. For each subject, we select two seed articles with more than 100 citations since 2020. For each article, we retrieve related articles from Semantic Scholar. [^8] Semantic Scholar provides APIs to retrieve scientific articles with metadata including titles, authors, a reference list with articles that the current article cites, and a citation list with articles that cite the current article. Throughout this paper, a reference is a bibliographic entry listed in the reference section of a scientific article, whereas a citation is an in-text acknowledgement of another’s work within the article. To ensure broad coverage, we include two-hop related articles identified via Semantic Scholar’s citation graph based on the reference and citation list. For each seed article, we randomly sample ten first-hop related articles, and for each of these, we further sample five second-hop related articles to form an article cluster. Since full texts are required, we filter out articles without PDFs. This process yields 16 article clusters comprising 662 scientific articles with PDFs across eight subjects (see step I in Figure 1). Finally, we convert the collected PDFs into markdown texts using Marker. [^9] https://github.com/datalab-to/marker
From these clusters, we create article collections of varying context lengths for our question-answering task. By concatenating the markdown texts, we generate collections with lengths of 64K, 128K, 512K, and 1M tokens. [^10] We can easily construct collections with arbitrary lengths exceeding 1M tokens. Each collection is initialized with a randomly selected article, and additional articles are added until it reaches a specified length level. We expand the collections using two strategies: (1) random sampling from the clusters, and (2) traversing the citation graph within each cluster (using both depth-first and breadth-first search) to construct collections that preserve citation relations among articles. Each collection contains at least four articles, and no two collections share more than half of their articles. From the 662 scientific articles across 16 clusters, we construct 2,612 article collections spanning all four length levels, including 2,027 generated through random sampling and 585 through graph traversal.
@@ 3.2 Creating Databases and SQL Queries
Once the article collections are assembled, we construct a database for each (see step II in Figure 1), focusing on a core subset of elements shared across all articles: titles, authors, and references. Based on these key elements, each database comprises three tables: articles, article-author, and citing-cited. The articles table contains metadata such as the title, reference count, and title word count. The article-author table captures author information, including names and their positions in the corresponding author list. The citing-cited table records citation relations among articles. A detailed description is provided in Appendix A. This information is obtained from Semantic Scholar or derived via simple preprocessing, such as splitting titles and counting words. Manual corrections were applied as needed to align with the full-text Markdown articles.
@@
We collectively refer to the SQL commands in Table 4 as information processing skills, since they test different information processing capabilities. We have various templates per skill, designed to be applicable across article collections. For each collection, we randomly select 10 templates and instantiate them with collection-specific values for all placeholders. We then execute these queries against the corresponding database to generate ground-truth answers.

### v3

# arXiv:2509.21028v3[cs.AI] 01 Mar 2026

@@ Abstract
We introduce SciTrek, a diagnostic question-answering benchmark designed to probe long-context numerical reasoning in large language models (LLMs). Existing long-context benchmarks mostly focus on simple information retrieval, rely on artificial contexts, or leave numerical reasoning unexplored. SciTrek addresses these limitations through questions that require counting, sorting, aggregating, and comparing information across multiple full-text scientific articles. Questions are automatically generated by formulating them as SQL queries over a database constructed from article metadata (titles, authors, and references), with ground-truth answers obtained via query execution. This design provides verifiable reasoning traces for fine-grained error analysis and enables efficient scaling to longer contexts with minimal human supervision. Extensive experiments on thirteen frontier open-weight and proprietary LLMs reveal that SciTrek poses a significant challenge: even the best-performing model achieves only 46.5% exact match at 128K tokens, with performance declining as the context length increases. Models particularly struggle with citation-related questions and compound logical conditions, including negation. [^1] Our dataset and code are available from https://github.com/oaimli/SciTrek
@@ 1 Introduction
Large language models (LLMs) now routinely support context windows of 128K tokens, with some exceeding one million (Gemini Team, 2025; OpenAI, 2025; Yang et al., 2025). Yet whether these models can actually reason over such extended inputs — rather than merely accommodate them — remains an open question. As these models are increasingly deployed in applications that demand information integration across lengthy documents (Liu et al., 2025), rigorous evaluation of their reasoning capabilities over long contexts becomes essential. A growing number of benchmarks have been developed to this end (Kuratov et al., 2024; Maekawa et al., 2025; Bai et al., 2025; Yen et al., 2025), but existing efforts suffer from three key limitations.
First, most focus on simple information retrieval rather than multi-step reasoning. Needle-In-A-Haystack tasks (Kamradt, 2023; Hsieh et al., 2024), for instance, have largely saturated for state-of-the-art long-context models (Gemini Team, 2025; OpenAI, 2025; Yang et al., 2025). Second, benchmarks that do require reasoning typically rely on artificial contexts created by injecting noise into short passages (Kuratov et al., 2024) (e.g., HotpotQA; Yang et al. 2018) or by synthesizing inputs with LLMs (Maekawa et al., 2025; Wang et al., 2024b), which limits ecological validity. Benchmarks built on natural long contexts, by contrast, depend on substantial human annotation, making them difficult to scale to rapidly expanding context windows and increasing the risk of data contamination and obsolescence (Asai et al., 2024; An et al., 2024; Wang et al., 2024c; Bai et al., 2025). Third, numerical reasoning, such as counting, aggregating, comparing, and sorting, remains largely unexplored in long-context evaluation, despite being fundamental to real-world tasks over documents (Wolfson et al., 2020, 2025).
In this paper, we introduce SciTrek, an analytical question-answering benchmark that addresses all three limitations. We generate questions requiring numerical operations (counting, aggregating, sorting, and comparing) over core elements of scientific articles: titles, authors, and references. Because these operations correspond to standard database functions, we formulate them as SQL queries over article metadata, obtaining ground-truth answers through query execution without manual annotation. We then exploit recent advances in LLM-based SQL-to-text generation (Yu et al., 2018; Li et al., 2023; Hui et al., 2024) to convert queries into natural-language questions, and validate that their meaning is preserved through round-trip verification (SQL $\to$ question $\to$ SQL).
As an example, consider the first question in Table 1: “What is the highest number of authors that any single article has?”. Although seemingly simple, answering it requires locating author lists across all articles in the context, counting authors per article, comparing counts, and identifying the maximum. These atomic reasoning operations underpin complex questions in expert-curated benchmarks (Wolfson et al., 2020, 2025) and recur in human-written questions in OpenScholar, LongBench v2, and Loong (Asai et al., 2024; Bai et al., 2025; Wang et al., 2024c) (see Appendix A for analysis). While our questions target surface-level metadata rather than domain-specific content, they test prerequisite capabilities: models that cannot reason accurately about easily identifiable elements are unlikely to succeed on deeper analytical tasks. Indeed, when models are given structured database tables rather than full texts and SQL queries are executed directly, accuracy rises sharply (Section 4). This feature confirms SciTrek’s diagnostic value: the gap between tabular and full-text performance isolates the specific cost of reasoning over unstructured long contexts, and the SQL traces provide interpretable reference solutions for fine-grained error analysis.
SciTrek contains a test set of 2,121 question-answer pairs on scientific articles and a training set of 19,543 question-answer pairs, with contexts of varying lengths (i.e., 64K, 128K, 512K, and 1M). [^2] In the paper, K refers to 1,024 and M to 1,024K. Table 1 shows examples of questions, their corresponding SQL queries, and answers. We use SciTrek to evaluate the numerical reasoning capabilities of frontier LLMs with context windows exceeding 128K tokens. This paper makes the following contributions:
  • We propose a methodology for constructing long-context numerical reasoning benchmarks from scientific articles, in which questions and verifiable answers are generated automatically via SQL queries over article metadata. The SQL queries additionally serve as interpretable reasoning traces, enabling fine-grained diagnosis of where and why models fail. The methodology requires minimal human intervention and generalises to any domain whose entities can be structured as database tables, making it straightforward to scale to longer contexts or larger datasets as model capabilities evolve.
  • Extensive experiments on thirteen frontier LLMs reveal that SciTrek poses a significant challenge: even the best model (o4-mini) achieves only 46.5% exact match at 128K tokens, with performance declining sharply as context length increases. All models struggle particularly with citation-related questions and sorting tasks. Supervised fine-tuning and reinforcement learning yield modest improvements, but do not close the gap.
  • Fine-grained error analysis exposes systematic failure patterns: models miscount bibliographic elements while performing better on smaller elements (title words and authors), misinterpret compound conditions, especially those involving negation, and frequently produce incomplete or incorrectly formatted outputs.
@@
Dataset Natural Context Numerical Scalable Length     Example Question
@@
OpenScholar 3K Compile a list of reviews […], and identify the most promising […]
LongBench v2 128K How long have I been living in my current apartment in Shinjuku?
@@
Loong 250K Which company has the highest non-current assets?
SciTrek 1M How many articles are cited by other articles but do not cite any other articles?
Table 2: Representative benchmarks for evaluating long-context language models. Natural Context: whether the benchmark uses naturally occurring input contexts; Numerical: whether the benchmark requires numerical reasoning over long contexts; Scalable: whether the benchmark can be extended with minimal human effort (e.g., to longer contexts or larger datasets); Length: maximum supported input length in tokens; : length is based on the retrieved text rather than the full articles which are not available.
@@ 2 Related Work
Retrieval-focused Benchmarks
@@
The increasing interest in long-context language models has driven the development of numerous benchmarks. Initial efforts primarily test whether models can locate specific information in long contexts, typically through Needle-In-A-Haystack (NIAH) tasks, such as NeedleBench (Li et al., 2024). Subsequent benchmarks extend this paradigm by adding distractor information or scattering relevant content across the input, including Ada-LEval (Wang et al., 2024a), BABILong (Kuratov et al., 2024), HELMET (Yen et al., 2025), LIFBENCH (Wu et al., 2025b), LongBench (Bai et al., 2024), and RULER (Hsieh et al., 2024). While these benchmarks usefully stress-test context retrieval, they do not evaluate whether models can reason over the retrieved information, particularly through numerical operations such as counting, aggregating, or comparing.
Reasoning over Synthetic Contexts
A second line of work targets reasoning over long contexts, but relies on artificial inputs. HoloBench (Maekawa et al., 2025) evaluates numerical reasoning over various types of tables, and MathHay (Wang et al., 2024b) focuses on mathematical reasoning within long documents. However, both benchmarks construct their contexts using model-generated content, which limits the realism of the evaluation and raises questions about whether the challenges they pose reflect those encountered in natural texts. Models may exploit distributional regularities in synthetic content that would not transfer to realistic documents.
Reasoning over Natural Contexts
Benchmarks that combine reasoning with natural contexts include LongBench v2 (Bai et al., 2025), LongMemEval (Wu et al., 2025a) and L-Eval (An et al., 2024). In addition, Loong Wang et al. (2024c), CURIE Cui et al. (2025), and OpenScholar Asai et al. (2024) evaluate reasoning over scholarly documents. However, these benchmarks rely on expert-curated questions and answers, which limits their scalability to longer contexts or larger datasets, increases the risk of data contamination as benchmark items are absorbed into model training data, and makes it difficult to adapt the benchmarks as model context windows expand. Moreover, because the reasoning steps required to arrive at correct answers are typically not provided, fine-grained failure analysis remains difficult. Among these, Loong is the closest to SciTrek in that it evaluates both numerical reasoning and uses natural contexts but it does not scale automatically and is limited to 250K tokens.
SciTrek takes a different approach. Rather than relying on human annotation, we generate questions automatically by constructing database tables over article metadata (titles, authors, references) and instantiating SQL query templates, which are then converted to natural language (using an LLM with round-trip verification). This approach draws on the broader tradition of SQL-based question generation Yu et al. (2018); Li et al. (2023); Hui et al. (2024). It ensures that every question has a verifiable ground-truth answer, the reasoning process is transparent through the underlying SQL query, and the benchmark can be scaled to arbitrary context lengths and new article collections with minimal human effort.
We compare SciTrek with existing benchmarks in Table 2. Scientific articles provide a particularly suitable testbed for long-context numerical reasoning: they contain rich relational structure through citations, authorship, and cross-references, and scale naturally to long contexts as article collections grow. The numerical operations that SciTrek evaluates are commonly required in human-annotated or human-audited benchmarks (see further analysis in Appendix A) and represent prerequisite capabilities for more complex analytical tasks.
@@ 3 The SciTrek Benchmark
Figure 1: Overview of SciTrek construction pipeline. We first gather scientific article collections of varying scales; then generate SQL queries and their answers from databases storing collection-specific metadata; and finally convert the SQL queries to natural language questions. The final dataset comprises three components highlighted in blue: full-text scientific articles as the input context, the natural language question, and the corresponding answer.
Figure 1 illustrates the SciTrek curation pipeline: we first assemble collections of scientific articles as contexts at varying lengths (i.e., 64K, 128K, 512K, and 1M), then construct databases representing article metadata and generate SQL queries with answers obtained via database execution. Finally, we convert the SQL queries into natural language questions. We describe each stage in detail below.
@@ 3.1 Gathering Scientific Article Collections
SciTrek is constructed from scientific articles sourced from Semantic Scholar. [^3] https://www.semanticscholar.org/ To ensure the representation of diverse topics, we obtain an initial set of seed articles from eight subjects: Computer Science (CS), Economics, Electronic Engineering (EE), Math, Physics, Biology, Finance, and Statistics. [^4] Following the typology of subjects from https://arxiv.org/. For each subject, we select two seed articles with more than 100 citations since 2020 and for each article, we retrieve related articles via the Semantic Scholar API. [^5] Semantic Scholar provides APIs to retrieve scientific articles together with metadata including titles, authors, a reference list with articles that the current article cites, and a citation list with articles that cite the current article. Throughout this paper, a reference is a bibliographic entry listed in the reference section of a scientific article, whereas a citation is an in-text acknowledgement of another’s work within the article. To ensure broad coverage, we include two-hop related articles identified via Semantic Scholar’s citation graph based on the reference and citation list. For each seed article, we randomly sample ten first-hop neighbors, and for each of these we further sample five second-hop neighbors froming an article cluster. Since full texts are required, we filter out articles without PDFs. This process yields 16 article clusters comprising 662 scientific articles with PDFs across eight subjects (see step I in Figure 1). Finally, we convert the collected PDFs to markdown using Marker. [^6] https://github.com/datalab-to/marker
From these clusters, we create article collections at varying context lengths for our question-answering task. By concatenating the Markdown texts, we produce collections of 64K, 128K, 512K, and 1M tokens. [^7] We can easily construct collections with arbitrary lengths exceeding 1M tokens. Each collection is initialised with a randomly selected article, and further articles are appended until the target length is reached. We employ two expansion strategies: (1) random sampling from the clusters, and (2) traversal of the citation graph within each cluster (using both depth-first and breadth-first search) to construct collections that preserve citation relations among articles. Random groupings expose models to heterogeneous, cross-domain contexts, while citation-based groupings introduce richer relational structure that demands more sophisticated long-context reasoning. Each collection contains at least four articles, and no two collections share more than half of their articles. From the 662 scientific articles across 16 clusters, we construct 2,612 collections spanning all four length levels: 2,027 via random sampling and 585 via graph traversal.
@@ 3.2 Creating Databases and SQL Queries
Once the article collections are assembled, we construct a database for each (see step II in Figure 1), focusing on a core subset of elements shared across all articles: titles, authors, and references. Based on these key elements, each database comprises three tables: articles, article-author, and citing-cited. The articles table contains metadata such as the title, reference count, and title word count. The article-author table captures author information, including names and their positions in the corresponding author list. The citing-cited table records citation relations among articles. A detailed description is provided in Appendix B. This information is obtained from Semantic Scholar or derived via simple preprocessing, such as splitting titles and counting words. Manual corrections were applied as needed to align with the full-text Markdown articles.
@@
We collectively refer to the SQL commands in Table 4 as information processing skills, since they test different information processing capabilities. We have various templates per skill, designed to be applicable across collections. For each collection, we randomly select 10 templates and instantiate them with collection-specific values for all placeholders. We then execute these queries against the corresponding database to generate ground-truth answers.